A rapidly spinning paddle wheel raises the temperature of 450 mL of water from 22.8ºC to 25.9ºC.  (Hint: to find mass, use the relation between, density, mass and volume) 

a) The mass of water in grams

(density of water = 1000 kg/m3, also convert volume of water to cubic centimeter 1mL=1cm3).

 

b) The difference in temperature in kelvin

 

b) The work (energy) done in this process in joules 

(specific heat of water = 4186 J/kg°C)

Grade:6

1 Answers

Gaurav
2 years ago
We know ,desinty of water is =1000kg/metercube =1gm/ml.   
So,mass of water = 1×450= 450gm.
Difference of tempreture in KELVIN = Difference of tempreture in degree
So, difference of tempreture in KELVIN = 25.9-22.8=3.1Kelvin 
Total woke done = specific heat × mass of in kg×tempreture difference
Total work= 4186×0.450×3.1=5839.47joule